# Mean estimation in the add-remove model of differential privacy

### 11 Dec 2023 v1

@@ Abstract
Differential privacy is often studied under two different models of neighboring datasets: the add-remove model and the swap model. While the swap model is used extensively in the academic literature, many practical libraries use the more conservative add-remove model. However, analysis under the add-remove model can be cumbersome, and obtaining results with tight constants requires some additional work. Here, we study the problem of one-dimensional mean estimation under the add-remove model of differential privacy. We propose a new algorithm and show that it is min-max optimal, that it has the correct constant in the leading term of the mean squared error, and that this constant is the same as the optimal algorithm in the swap model. Our results show that, for mean estimation, the add-remove and swap model give nearly identical error even though the add-remove model cannot treat the size of the dataset as public information. In addition, we demonstrate empirically that our proposed algorithm yields a factor of two improvement in mean squared error over algorithms often used in practice.
@@ 1 Introduction
Mean estimation is one of the simplest and most widely used techniques in statistics, and is often deployed as a subroutine in more complex algorithms. However, the mean of a dataset can potentially reveal private information, and hence several differentially private methods to estimate it have been proposed. These include private mean estimation with robustness (Liu et al., 2021), mean estimation under statistical models (Kamath et al., 2020), and instance-specific results (Huang et al., 2021; Dick et al., 2023). Given the ubiquitous nature of differentially private mean estimation, we revisit the problem for scalar inputs and propose an algorithm that is min-max optimal up to the leading term (including the constant).
Before we proceed further, we state a few notations and definitions. Let $\mathcal{D}_{n}(\ell,u)$ denote the set of all datasets consisting of $n$ real values in the range $[\ell,u]$, let $\mathcal{D}_{\geq n}(\ell,u)$ denote the set of all datasets consisting of at least $n$ real values in the range $[\ell,u]$, and let $\mathcal{D}^{*}(\ell,u)$ denote the set of all datasets of all sizes. Differentially private mechanisms are defined as follows:
@@
A randomized, real-valued algorithm $A$ satisfies $\varepsilon$-differential privacy if for any two neighboring datasets $D,D^{\prime}$ in $\mathcal{D}^{*}(\ell,u)$ and for any output $\mathcal{S}\subseteq\mathcal{R}$, it holds that
@@
Two definitions of neighboring datasets are popular. The first definition is called the swap model of neighboring datasets (Dwork et al., 2006; Vadhan, 2017), where two datasets $D$ and $D^{\prime}$ are neighboring if and only if
@@
The second definition is called the add-remove model of neighboring datasets (Dwork et al., 2014, Definition 2.4), where two datasets $D$ and $D^{\prime}$ are neighboring if and only if
@@
While both neighborhood models of differential privacy are studied in the literature, the add-remove model is often used for statistical queries in practice (McSherry, 2009; Wilson et al., 2019; Rogers et al., 2020; Amin et al., 2022). This may be because it protects the size of the dataset, while the swap model does not. Furthermore, any $\varepsilon$-DP algorithm under the add-remove model is also an $2\varepsilon$-DP algorithm under the swap model, so in this sense the add-remove model is more conservative than the swap model.
Let $\mathcal{A}^{\text{ar}}_{\varepsilon}$ denote the set of all $\varepsilon$-differentially private algorithms under the add-remove model and $\mathcal{A}^{\text{sw}}_{\varepsilon}$ denote the set of all $\varepsilon$-differentially private algorithms under the swap model. Given a dataset $D=\{x_{1},x_{2},\ldots,\}$, the mean of the dataset is given by
@@
We measure the utility of a mean estimator $\hat{\mu}:\mathcal{D}^{*}(\ell,u)\to[\ell,u]$ for a dataset $D$ in terms of mean squared error,
@@
where the expectation is over the randomization of the algorithm. Typically, $L(\hat{\mu},D)$ decreases with the size of $D$ as $O(1/|D|^{2})$. Hence, we measure the normalized mean squared error as
$L_{\text{norm}}(\hat{\mu},D)\triangleq|\mathcal{D}|^{2}L(\hat{\mu},D),$
and the min-max normalized mean squared error for the add-remove model is defined as
@@
We can similarly define the min-max normalized mean squared error under the swap model as
$R_{\text{sw}}(\varepsilon,n_{0},\ell,u)\triangleq\inf_{\hat{\mu}\in\mathcal{A}^{\text{sw}}_{\varepsilon}}\sup_{D\in\mathcal{D}_{\geq n_{0}}(\ell,u)}L_{\text{norm}}(\hat{\mu},D).$
@@
$R_{\text{sw}}(\varepsilon,n_{0},\ell,u)=\frac{2(u-\ell)^{2}}{\varepsilon^{2}}\left(1\pm o(1)\right).$ (1)
Furthermore, the Laplace mechanism and the staircase mechanism achieve this mean-squared loss up to the $o(1)$ term (Geng and Viswanath, 2014). Throughout this paper, we use $o(1)$ to indicate a term that tends to zero as $n_{0}\to\infty$ and $\varepsilon\to 0$ such that $n_{0}\varepsilon\to\infty$. For simplicity, we drop the term $n_{0}$ in the notation of $R_{\text{sw}}$ and $R_{\text{ar}}$ in the rest of the paper.
The story is a bit more complicated under the add-remove model. Recall that the mean is the ratio of the sum ($s\triangleq\sum_{x\in D}x$) and the count ($n\triangleq|D|$). A simple algorithm to estimate the mean is to use a fraction of the privacy budget (say, $\varepsilon/2$) to estimate the sum $\hat{s}$, use the remaining privacy budget ($\varepsilon/2$) to estimate the count $\hat{n}$, and finally estimate the mean by taking the ratio $\hat{s}/\hat{n}$. Since the true mean always lies in the range $[\ell,u]$, we can further clip the estimated ratio to $[\ell,u]$ to improve accuracy, where the clipping function is given as follows:
$\text{Clip}(x,[a,b])=\max(a,\min(x,b)).$
This standard algorithm is given in Algorithm 1. The noise added in Algorithm 1 is proportional to $\max(|\ell|,|u|)$, which can be prohibitive for many datasets.
@@
Algorithm 2 is a minor modification to Algorithm 1 where the inputs are shifted by $(\ell+u)/2$ before computing the sum. The shift reduces the sensitivity of the sum to $(u-\ell)/2$, which improves the error. It can be shown that under this algorithm, for any dataset $D$,
@@
and hence,
$R_{\text{ar}}(\varepsilon,\ell,u)\leq\frac{4(u-\ell)^{2}}{\varepsilon^{2}}\left(1+o(1)\right).$ (2)
The factor two difference between the bounds on add-remove model (2) and the swap model (1) raises a natural question: is the mean estimation under the add-remove model inherently harder? We answer this question in this paper.
@@ 2 Our contributions
For the sake of completeness, we first analyze Algorithm 2 and provide a dataset specific upper bound in Lemma 2 by proving that for any dataset $D$,
$\mathbb{E}[(\hat{\mu}(D)-\mu(D))^{2}]$ $\leq\left(\frac{2(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{8(\mu-\frac{u+\ell}{2})^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right).$
We then propose a new mean estimation algorithm in Algorithm 3. Instead of computing the sum and count as in Algorithm 2, the new algorithm computes the scaled sum (denoted by $s_{1}$) and the difference of count and scaled sum (denoted by $s_{2}$). It then privatizes both $s_{1}$ and $s_{2}$ and uses it to compute mean. In contrast to Algorithm 2, Algorithm 3 adds correlated noise to the numerator and denominator. Using this observation, in Theorem 1, we show that the new algorithm has error at most
$\mathbb{E}[(\hat{\mu}(D)-\mu(D))^{2}]$ $\leq\left(\frac{(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{4(\mu-\frac{u+\ell}{2})^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right).$
@@
Furthermore, in Theorem 2, we show an information-theoretic lower bound:
$R_{\text{ar}}(\varepsilon,\ell,u)\geq\frac{2(u-\ell)^{2}}{\varepsilon^{2}}\left(1-o(1)\right),$
thus establishing that
$R_{\text{ar}}(\varepsilon,\ell,u)=\frac{2(u-\ell)^{2}}{\varepsilon^{2}}\left(1\pm o(1)\right).$
Note that the leading order terms in the upper bounds on MSE for Algorithms 2 and 3 differ by a factor of two. In Section 6, we show that the ratio of true MSE between the two algorithms is two for several datasets.
We note that while Algorithm 3 uses Laplace mechanism as a subroutine, it can be substituted by the staircase mechanism (Geng and Viswanath, 2014). Furthermore, when the bounds $\ell$ and $u$ are unknown and data is possible unbounded, the above algorithm can be combined with standard clipping algorithms to get algorithms that perform on unbounded domains (Amin et al., 2019).
@@
Let $\hat{s}_{1}=s_{1}+Z_{1}$, where $Z_{1}\sim\text{Lap}(\frac{1}{\varepsilon})$.
Let $\hat{s}_{2}=s_{2}+Z_{2}$, where $Z_{2}\sim\text{Lap}(\frac{1}{\varepsilon})$.
@@
The rest of the paper is organized as follows. In Section 3, we motivate our approach by showing that applying a linear transformation to the [sum, count] vector could lead to a better algorithm. In Section 4, we analyze the algorithms and provide upper bounds on their mean squared error. In Section 5, we provide an information-theoretic lower bound. In Section 6, we empirically demonstrate the performance of the new algorithm.
3 Improving the Laplace mechanism via linear transformation
We first motivate the approach via K-Norm mechanisms (Hardt and Talwar, 2010). The Laplace mechanism belongs to the more general family of K-Norm mechanisms, where a zero-mean noise is added, whose probability distribution at any given point exponentially decreases in proportion to a defined norm. We first view Algorithm 2 in this perspective.
Without loss of generality, we assume $\ell=0$ and $u=1$. Let $f(S)=[\sum_{x\in S}x,\sum_{x\in S}1]$ be the two-dimensional vector output for multiset $S$. The sensitivity space is defined by all possibilities of $f(S)-f(S^{\prime})$ for neighboring $S$ and $S^{\prime}$. In this context, it is ±$\pm[x,1]$ for $x\in[0,1]$, as shown by the two bold segments in the 2D picture below in the middle of Figure 1. K-norm mechanism defines a norm $\|\|_{K}$, whose unit ball is a convex shape $K$. The probability of sampling a noise $P(x)\propto\exp(-\varepsilon\|x\|_{K})$. If $K$ contains the whole sensitivity space, this mechanism is guaranteed to preserve $\varepsilon$-DP.
For the Laplace mechanism, the unit ball is $\{x:\|x\|_{1}=a\}$ for some constant $a$. The smallest such ball that could cover the whole sensitivity space has $a=2$, given as the red square in the middle of Figure 1. Noticeably, it wastefully covers way more area than the convex hull of sensitivity space: the parallelogram ${(0,1),(1,1),(0,-1),(-1,-1)}$ (in blue).
Compared to this naive algorithm, Algorithm 2 is better, since the shift in Line 2 brings all entries from range $[0,1]$ to $[-\frac{1}{2},\frac{1}{2}]$, which helps reduce the sensitivity and hence the scale of Laplace noise. In fact, one can show that Algorithm 2 is equivalent to the following procedure:

### 19 Feb 2024 v2

@@ Abstract
Differential privacy is often studied under two different models of neighboring datasets: the add-remove model and the swap model. While the swap model is frequently used in the academic literature to simplify analysis, many practical applications rely on the more conservative add-remove model, where obtaining tight results can be difficult. Here, we study the problem of one-dimensional mean estimation under the add-remove model. We propose a new algorithm and show that it is min-max optimal, achieving the best possible constant in the leading term of the mean squared error for all $\varepsilon$, and that this constant is the same as the optimal algorithm under the swap model. These results show that the add-remove and swap models give nearly identical errors for mean estimation, even though the add-remove model cannot treat the size of the dataset as public information. We also demonstrate empirically that our proposed algorithm yields at least a factor of two improvement in mean squared error over algorithms frequently used in practice. One of our main technical contributions is a new hourglass  mechanism, which might be of independent interest in other scenarios.
@@ 1 Introduction
Mean estimation is one of the simplest and most widely used techniques in statistics, and is often deployed as a subroutine for more complex analyses. However, the mean of a dataset can reveal private information, and so a variety of differentially private methods have been proposed to estimate it. These include private mean estimation with robustness (Liu et al., 2021), mean estimation under statistical models (Kamath et al., 2020), and techniques with instance-specific guarantees (Huang et al., 2021; Dick et al., 2023).
However, despite the ubiquitous nature of differentially private mean estimation, fundamental questions remain about the optimal error that can be achieved under commonly used variants of differential privacy. We begin with the general definition of a differentially private mechanism.
@@
A randomized, real-valued algorithm $A$ satisfies $\varepsilon$-differential privacy if for any two neighboring datasets $D,D^{\prime}$ and for any output $\mathcal{S}\subseteq\mathcal{R}$, it holds that
@@
It remains to specify what makes two datasets neighbors. Two definitions are commonly used. The first, called the swap model (Dwork et al., 2006; Vadhan, 2017), defines two datasets $D$ and $D^{\prime}$ as neighboring if and only if
@@
The second, called the add-remove model (Dwork et al., 2014, Definition 2.4), defines $D$ and $D^{\prime}$ as neighboring if and only if
@@
Intuitively, under the swap model, $D^{\prime}$ is obtained from $D$ by changing a value, while under the add-remove model, it is obtained by adding or removing a value.
While both neighborhood models of differential privacy are studied in the literature, the add-remove model is more frequently used for statistical queries in practice (McSherry, 2009; Wilson et al., 2019; Rogers et al., 2020; Amin et al., 2022), likely because it is more conservative: the add-remove model protects the size of the input dataset, while the swap model does not. Furthermore, a $\varepsilon$-DP algorithm under the add-remove model is also a $2\varepsilon$-DP algorithm under the swap model, so in this sense the relationship is strict.
Here, we revisit the problem of scalar mean estimation in the add-remove model of differential privacy, proposing a new algorithm that is min-max optimal, including the constant on the leading term of the error, and showing that the add-remove and swap models give nearly identical errors despite the former’s additional protections.
1.1 Setting
Let $\mathcal{D}_{n}(\ell,u)$ denote the set of all datasets consisting of $n$ real values in the range $[\ell,u]$, let $\mathcal{D}_{\geq n}(\ell,u)$ denote the set of all datasets consisting of at least $n$ real values in the range $[\ell,u]$, and let $\mathcal{D}^{*}(\ell,u)$ denote the set of datasets of all sizes with values in $[\ell,u]$.
Given a dataset $D=\{x_{1},x_{2},\ldots,\}$, the mean is given by
@@ 1.1 Setting
We measure the utility of a mean estimator $\hat{\mu}:\mathcal{D}^{*}(\ell,u)\to[\ell,u]$ for a dataset $D$ in terms of mean squared error (MSE),
@@
where the expectation is over the randomization of the estimator. Typically, for private estimators, $L(\hat{\mu},D)$ decreases with the size of $D$ as $O(1/|D|^{2})$. Hence, we measure the normalized mean squared error as
$L_{\text{norm}}(\hat{\mu},D)\triangleq|D|^{2}L(\hat{\mu},D).$
Let $\mathcal{A}^{\text{sw}}_{\varepsilon}$ denote the set of all $\varepsilon$-differentially private algorithms under the swap model, and let $\mathcal{A}^{\text{ar}}_{\varepsilon}$ denote the set of all $\varepsilon$-differentially private algorithms under the add-remove model. The min-max normalized mean squared error for sufficiently large datasets in the swap model is defined as
$R_{\text{sw}}(\varepsilon,n_{0},\ell,u)\triangleq\inf_{\hat{\mu}\in\mathcal{A}^{\text{sw}}_{\varepsilon}}\sup_{D\in\mathcal{D}_{\geq n_{0}}(\ell,u)}L_{\text{norm}}(\hat{\mu},D),$
and similarly in the add-remove model is
@@
1.2 Optimality in the swap model
@@
$R_{\text{sw}}(\varepsilon,n_{0},\ell,u)=(u-\ell)^{2}\cdot\sigma^{2}(\varepsilon),$ (1)
where
$\sigma^{2}(\varepsilon)=\frac{2^{-2/3}e^{-2\varepsilon/3}(1+e^{-\varepsilon})^{2/3}+e^{-\varepsilon}}{(1-e^{-\varepsilon})^{2}}.$ (2)
As $\varepsilon\to 0$, $\sigma^{2}(\varepsilon)\to 2/\varepsilon^{2}$, and hence for small values of $\varepsilon$
$R_{\text{sw}}(\varepsilon,n_{0},\ell,u)=\frac{2(u-\ell)^{2}}{\varepsilon^{2}}\left(1\pm o(1)\right),$ (3)
where the $o(1)$ term tends to zero as $n_{0}\to\infty$. The Laplace mechanism matches this mean squared error up to the $o(1)$ term, and in this sense is optimal as $\varepsilon\to 0$.
However, for larger $\varepsilon$, the Laplace mechanism is not optimal. Geng and Viswanath (2014) defined a class of differentially private mechanisms called staircase mechanisms whose density is parameterized by $\gamma\in[0,1]$ and showed that, for any monotonic loss, there exists a $\gamma$ such that the staircase mechanism is the optimal differentially private mechanism for that loss. We provide a definition of the staircase mechanism in Definition 2 for completeness.
1.3 Optimality in the add-remove model
The story is a bit more complicated under the add-remove model. Since the mean is the ratio of the sum ($s\triangleq\sum_{x\in D}x$) to the count ($n\triangleq|D|$), one simple algorithm for private mean estimation is to use a fraction of the privacy budget (say, $\varepsilon/2$) to estimate the sum as $\hat{s}$, use the remaining privacy budget ($\varepsilon/2$) to estimate the count as $\hat{n}$, and finally estimate the mean as $\hat{s}/\hat{n}$. Since the true mean always lies in the range $[\ell,u]$, we can additionally clip the result to $[\ell,u]$ to improve accuracy. This standard algorithm is shown in Algorithm 1, where $\text{Clip}(x,[a,b])=\max(a,\min(x,b))$.
@@ 1.3 Optimality in the add-remove model
The noise added in Algorithm 1 is proportional to $\max(|\ell|,|u|)$, which can be badly suboptimal, for instance if $l=10^{6}$ and $u=10^{6}+1$. Algorithm 2 modifies Algorithm 1 by shifting the inputs by $(\ell+u)/2$ before computing the sum, reducing the sensitivity to $(u-\ell)/2$. This improves the error, sometimes dramatically. It can be shown that, for any dataset $D\in\mathcal{D}^{*}(\ell,u)$,
@@
and hence
$R_{\text{ar}}(\varepsilon,n_{0},\ell,u)\leq\frac{4(u-\ell)^{2}}{\varepsilon^{2}}\left(1+o(1)\right).$ (4)
As before, $o(1)$ indicates a term that vanishes as the minimum dataset size $n_{0}$ grows; for simplicity, we drop $n_{0}$ in the notation of $R_{\text{sw}}$ and $R_{\text{ar}}$ for the rest of the paper.
This result is still a factor of two larger than the swap model lower bound in (3), and the loss is even further from $R_{\text{sw}}$ in the low-privacy regime where $\varepsilon$ is large. Recently, Kamath et al. (2023) proposed a generic mechanism to convert a differentially private algorithm in the swap model to a differentially private algorithm in the add-remove model and instantiated it for the unbiased mean estimation problem. However, their focus was not the constant in the mean squared error, and the stated result (Kamath et al., 2023, Theorem D.6) has a constant of $9$.
Thus, a natural question remains: is mean estimation in the add-remove model inherently harder than in the swap model? We show in this paper that the answer is no.
@@ 2 Our contributions
We propose a new mean estimation algorithm in Algorithm 3, introducing two key improvements over Algorithm 2:
Transformed noise addition. Instead of estimating the sum and count directly as in Algorithm 2, the new algorithm estimates a linear transformation of the sum and count to reduce $\ell_{1}$ sensitivity. This is sufficient to achieve optimal error using the vector Laplace mechanism in the high-privacy regime where $\varepsilon$ is small.
The hourglass mechanism. To achieve optimal error in the low-privacy regime where $\varepsilon$ is large, we propose a new two-dimensional noise distribution called the hourglass  mechanism. It has the desirable property that the marginal distribution over either dimension is the optimal univariate staircase mechanism (Geng and Viswanath, 2014).
We show that the hourglass mechanism can be sampled efficiently, and prove a bound on the mean squared error of Algorithm 3 when the noise is drawn from the hourglass mechanism. Combined with an information-theoretic lower bound on $R_{\text{ar}}(\varepsilon,\ell,u)$, this shows that Algorithm 3 is optimal for all $\varepsilon$, $\ell$, and $u$.
These bounds also match the result of Geng and Viswanath (2014) for $R_{\text{sw}}(\varepsilon,\ell,u)$, establishing that the swap model and the add-remove model give the same mean squared error (up to $o(1)$ terms).
We note in passing that, when the bounds $\ell$ and $u$ are unknown, Algorithm 3 can be combined with standard clipping algorithms to perform on unbounded domains (Amin et al., 2019).
2.1 Overview of technical results
We first analyze Algorithm 2 as a baseline and provide a dataset specific upper bound in Lemma 2, proving that for any dataset $D$, its mean squared error is upper bounded by
$\left(\frac{2(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{8(\mu-\frac{u+\ell}{2})^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right).$
We then analyze Algorithm 3 with Laplace noise in Theorem 1, showing that its error is at most
$\left(\frac{(u-l)^{2}}{|D|^{2}\varepsilon^{2}}+\frac{4(\mu-\frac{u+\ell}{2})^{2}}{|D|^{2}\varepsilon^{2}}\right)\left(1+o(1)\right),$ (5)
@@ 2.1 Overview of technical results
This is the same as the min-max MSE of the swap model for small values of $\varepsilon$.
We next analyze Algorithm 3 with noise drawn from the two-dimensional staircase mechanism proposed by Geng et al. (2015) in Lemma 5, and show that its error is at most
$\left(\frac{(u-l)^{2}\widetilde{\sigma}^{2}(\varepsilon)}{|D|^{2}}\right)\left(1+o(1)\right),$
where $\widetilde{\sigma}^{2}(\varepsilon)$ is the variance of the two-dimensional staircase mechanism optimized for mean squared error with privacy guarantee $\varepsilon$. While the above result is better than (5) for large values of $\varepsilon$, it still does not match the error of the swap model in general. This is due to the fact that for large values of $\varepsilon$, the error of the swap model scales as
$\sigma^{2}(\varepsilon)=\Theta\left((u-\ell)^{2}e^{-2\varepsilon/3}\right),$
while the error of the two dimensional staircase mechanism applied in Algorithm 3 scales as
$\widetilde{\sigma}^{2}(\varepsilon)=\Theta\left((u-\ell)^{2}e^{-\varepsilon/2}\right).$
(See Lemma 7 in the Appendix for details.)
We finally analyze Algorithm 3 with noise drawn from the hourglass mechanism in Theorem 3, showing that its error is upper bounded by
$\left(\frac{(u-l)^{2}{\sigma}^{2}(\varepsilon)}{|D|^{2}}\right)\left(1+o(1)\right),$
where $\sigma^{2}(\varepsilon)$ is given in (2), matching the swap model lower bound in (1).
In addition, we prove an information-theoretic lower bound for the add-remove model in Theorem 4:
$R_{\text{ar}}(\varepsilon,\ell,u)\geq(u-\ell)^{2}\cdot\sigma^{2}(\varepsilon)\cdot\left(1-o(1)\right),$
establishing that
$R_{\text{ar}}(\varepsilon,\ell,u)=(u-\ell)^{2}\cdot\sigma^{2}(\varepsilon)\cdot\left(1\pm o(1)\right),$
and therefore that Algorithm 3 is optimal, as well as that the add-remove and swap models have equivalent mean squared error for mean estimation.
@@
Let $Z=[Z_{1},Z_{2}]\sim\text{two-dim. noise mechanism}(\varepsilon)$.
Let $\hat{s}_{1}=s_{1}+Z_{1}$.
Let $\hat{s}_{2}=s_{2}+Z_{2}$.
@@
The rest of the paper is organized as follows. In Section 3 we discuss the high-privacy regime, showing how a linear transformation on the sum and count leads to optimal error using the Laplace mechanism. In Section 4, we generalize our results result to the low-privacy regime, introducing the hourglass mechanism in Section 4.1 and applying it to mean estimation in Section 4.2. In Section 5 we prove an information-theoretic lower bound showing that our results with the hourglass mechanism are optimal in the add-remove model and match the optimal error in the swap model as well. Finally, in Section 6, we empirically demonstrate the performance of our algorithm.
Figure 1: The Laplace mechanism applied to different linear transformations of $S(q)$. The middle plot shows the original sensitivity space, where red denotes the noise ball used by Algorithm 1; orange the ball used by Algorithm 2, and blue the ball used by Algorithm 3, which is the smallest convex shape possible. The left plot shows the transformed space used by Algorithm 3, and the right plot shows the transformed space used by Algorithm 2.
3 High privacy regime
We first build intuition by viewing Algorithms 1 and 2 geometrically, drawing on the framework of Hardt and Talwar (2010). To simplify, we will assume $\ell=0$ and $u=1$. Let
$q(D)=\left[\sum_{x\in D}x,\sum_{x\in D}1\right]$
be the two-dimensional sum and count vector for dataset $D$. Define the sensitivity space $S(q)$ to be the set of possible values for $q(D)-q(D^{\prime})$ when $D$ and $D^{\prime}$ are neighboring datasets. Under the add-remove model, the sensitivity space for $q$ is ±$\pm[x,1]$ for $x\in[0,1]$, depicted by the two bold line segments in the middle plot of Figure 1.
The standard vector Laplace mechanism can be used to obtain a differentially private estimate of $q(D)$ based on its sensitivity. In particular, the mechanism adds noise scaled to the maximum $\ell_{1}$ norm of the sensitivity space—that is, the smallest constant $a$ such that the $\ell_{1}$ ball $\{x:\|x\|_{1}=a\}$ contains $S(q)$. Here, the minimum value of $a$ is 2, as shown by the red diamond in the middle plot of Figure 1. And, indeed, adding Laplace noise scaled to $a=2$ is precisely what Algorithm 1 does.
However, noise added in this way actually supports a much larger sensitivity space than just $S(q)$, as can be seen in the figure. This means that Algorithm 1 is effectively “wasting” noise to protect against changes to $q$ that cannot occur, unnecessarily increasing error. (The problem is even worse when the range $[\ell,u]$ is far from zero, as noted earlier.)
Compared to this naive approach, Algorithm 2 is significantly better, since the shift in Line 2 maps values in the dataset from $[0,1]$ to $[-\frac{1}{2},\frac{1}{2}]$. The bold line segments in the right plot of Figure 1 depict the new sensitivity space, which is contained by the $\ell_{1}$ ball with smaller scale $a=1$ (shown in orange).
In fact, one can show that Algorithm 2 is equivalent to the following procedure: